# Geometric Distortion Guided Transformer for Omnidirectional Image Super-Resolution

### v1

@@ IV-C Ablation Study
Impacts of Diff. In DFA, we calculate the difference between $F^{DMRSA}$ and $F^{DDSA}$, which is used to adaptively weigh the combined features. As shown in Table III, the performance slightly degrades without Diff. Fig. 12 shows the visual comparison between our model without and with the Diff design in DFA. We can see that some details are missing without Diff, while more details are restored with Diff. It is worth noting that computing the Diff does not require additional parameters, but it can bring improvements in restoring details. In Table III, we can observe that the SR performance is moderately impacted when both DGG and Diff are absent simultaneously, especially under the ×$\times 2$ scaling factor. This demonstrates the necessity and importance of both utilizing the distortion map and leveraging the distinction between features from DMRSA and DDSA.
Impacts of Dynamic Feature Aggregation. We propose to aggregate features from DMRSA and DDSA in a dynamic way, based on the learned weights. To verify the efficacy of DFA, we compare the SR performance of different aggregation methods on the SUN 360 panorama dataset, as illustrated in Table V. ‘Addition’ denotes that the two features are summed directly to obtain the final feature. ‘Concatenation’ denotes that the two features are concatenated along the channel dimension, and then convolved by a convolutional layer to reduce the number of channels to match that of the input features. From Table V, we can see that DFA achieves the best or comparative performance across all evaluation metrics compared with other aggregation methods.
Impacts of Window Size. We investigate the effects of different sizes of rectangle windows, as shown in Table VI. A window size of (32/8) means that the height and width are 32 and 8 for vertical windows, respectively, and 8 and 32 for horizontal windows. As demonstrated in Table VI, the network with a window size of (64/8) outperforms those with window sizes of (32/8) and (64/4). This improvement may be because the rectangle window with size (64/8) can effectively capture features of distorted content in a more appropriate range of area.
@@
Impacts of Training Loss. Given the non-uniform pixel distribution, we adopt the WS-$l1$ loss as our training loss function to calculate pixel-wise errors in ERP images. From Table VII, we can see that WS-$l1$ outperforms $l1$ in the evaluation metrics of PSNR and WS-PSNR, while $l1$ performs better than WS-$l1$ in the evaluation metrics of SSIM and WS-SSIM. Referred to Eq. (23), it is probable that in the WS-$l1$ loss, luminance and contrast are affected by the multiplication of the distortion map in the training process, which leads to the degradation of SSIM and WS-SSIM. However, the performance of PSNR and WS-PSNR demonstrates that WS-$l1$ loss can improve and relieve pixel-wise errors.
@@
Exploration on Involved Area. To investigate the effects of the range of involved pixels, we randomly selected four different pixels from each image in the SUN 360 Panorama testing dataset and calculated their average DI values, as shown in Table VIII. Higher DI values indicate a wider range of pixels involved. Note that the coordinates of the upper left corner of an image are (0,0). As shown in Table VIII, the proposed method achieves the highest DI values at the four selected pixels from different latitudes compared to other methods, which demonstrates that our method has the widest range of involved area. Fig. 1 visually shows that our method’s area of contribution for reconstructing the patch in the red box is the widest with more self-similar textures, resulting in superior SR performance. These results indicate that the proposed mechanisms can enlarge the attention area and improve SR performance.
V Conclusion
@@
In this paper, we propose a Geometric Distortion Guided Transformer for Omnidirectional image Super-Resolution, named GDGT-OSR. Considering the geometric distortion in ERP images, we propose a Distortion Modulated Rectangle-window Self-Attention (DMRSA) mechanism, integrated with Distortion-aware Deformable Self-Attention (DDSA), to adapt to the unevenly distorted content. In this way, GDGT-OSR captures features from the attention areas with various shapes, aiming to calibrate the attention regions and facilitate their expansion, capturing more self-similar and related textures. To exploit the distortion map, we propose a Distortion Guidance Generator (DGG) to transform geometric distortion into distortion guidance, which is leveraged to modulate the key and value features in Rwin-SA. Furthermore, we adaptively aggregate two features generated by DMRSA and DDSA through a Dynamic Feature Aggregation (DFA) module. Experiment results demonstrate that GDGT-OSR can restore more details and richer textures over other methods, achieving state-of-the-art performance on omnidirectional image super-resolution.

### v2

@@ IV-C Ablation Study
Impacts of Training Loss. Given the non-uniform pixel distribution, we adopt the WS-$l1$ loss as our training loss function to calculate pixel-wise errors in ERP images. From Table IX, we can see that WS-$l1$ outperforms $l1$ in the evaluation metrics of PSNR and WS-PSNR, while $l1$ performs better than WS-$l1$ in the evaluation metrics of SSIM and WS-SSIM. Referred to Eq. (23), it is probable that in the WS-$l1$ loss, luminance and contrast are affected by the multiplication of the distortion map in the training process, which leads to the degradation of SSIM and WS-SSIM. However, the performance of PSNR and WS-PSNR demonstrates that WS-$l1$ loss can improve and relieve pixel-wise errors.
Eq. (23) indicates that the WS-$l1$ loss is obtained by weighing the $l1$ loss with the distortion map $D$. To further study the effectiveness and rationality of WS-$l1$ loss as the training loss, we train the model with a loss that weighs the $l1$ loss with (1-$D$). This loss assigns larger weights to the high-latitude areas and smaller weights to the low-latitude areas, which is opposite to the WS-$l1$ loss. It can easily be written as $l1-(\text{WS-}l1)$. As shown in Table IX, the model trained with $l1-(\text{WS-}l1)$ yields inferior results compared to those trained with $l1$ and WS-$l1$. These findings suggest that prioritizing lower-latitude regions, where pixels are more densely distributed, is crucial during the training process of ODISR.
IV-D Model Complexity
We compare the model complexity of our methods with other SOTA methods, as summarized in Table X. We evaluate the model performance on the ODI-SR dataset under the ×$\times 8$ scaling factor. The number of Multi-Adds is calculated for an input size of $64\times 64$. As shown in Table X, the number of parameters and Multi-Adds of HAT[31] are much larger than those of other methods. However, despite its substantial computational complexity, the performance of HAT[31] remains inferior to both OSRT[11] and our GDGT-OSR. On the contrary, compared to other methods, GDGT-OSR achieves significant performance improvements with a comparable or only marginally increased number of parameters. This highlights the superiority and effectiveness of our GDGT-OSR.
IV-E Exploration on Involved Area.
@@ IV-D Model Complexity
To investigate the effects of the range of involved pixels, we randomly selected four different pixels from each image in the SUN 360 Panorama testing dataset and calculated their average DI values, as shown in Table XI. Higher DI values indicate a wider range of pixels involved. Note that the coordinates of the upper left corner of an image are (0,0). As shown in Table XI, the proposed method achieves the highest DI values at the four selected pixels from different latitudes compared to other methods, which demonstrates that our method has the widest range of involved area. Fig. 1 visually shows that our method’s area of contribution for reconstructing the patch in the red box is the widest with more self-similar textures, resulting in superior SR performance. More visual results of local attribution maps are illustrated in Fig. 14, demonstrating that the proposed GDGT-OSR method effectively leverages a broader range of pixels. This capability enables it to produce more visually pleasing and higher-quality SR results. The above results indicate that the proposed mechanisms can enlarge the attention area and improve SR performance.
V Summary
V-A Limitations
Inference Speed. Typically, ODI images possess high resolutions, such as 2K, 4K, or even 8K. However, the inference speed of the proposed method is limited when processing these high-resolution ODI images due to the computationally intensive self-attention mechanisms employed. Furthermore, during inference, if the model size is excessively large, GPU memory may become insufficient to handle such high-resolution inputs. Although GDGT-OSR can produce high-quality SR results, its practical application is constrained in scenarios requiring real-time super-resolution of high-resolution ODI images with limited computational resources.
Figure 15: Failed cases under the ×$\times 16$ scaling factor.
Large Scaling Factors. When super-resolving objects with numerous high-frequency details, e.g., trees and humans, under the large scaling factor of ×$\times 16$, the qualitative results produced by GDGT-OSR are not visually satisfactory, as illustrated in Fig. 15. The primary reason is the significant loss of fine-grained information in the LR images when using a large scaling factor. With limited available information in the LR images, it becomes challenging for GDGT-OSR, as well as other SOTA methods, such as RGT[32] and OSRT[11], to reconstruct visually pleasing results for these high-frequency components in such demanding conditions.
V-B Conclusion
@@
In this paper, we propose a Geometric Distortion Guided Transformer for Omnidirectional image Super-Resolution, named GDGT-OSR. Considering the geometric distortion in ERP images, we propose a Distortion Modulated Rectangle-window Self-Attention (DMRSA) mechanism, integrated with Distortion-aware Deformable Self-Attention (DDSA), to adapt to the unevenly distorted content. In this way, GDGT-OSR captures features from the attention areas with various shapes, aiming to calibrate the attention regions and facilitate their expansion, capturing more self-similar and related textures. To exploit the distortion map, we propose a Distortion Guidance Generator (DGG) to transform geometric distortion into distortion guidance, which is leveraged to modulate the key and value features in Rwin-SA. Furthermore, we adaptively aggregate two features generated by DMRSA and DDSA through a Dynamic Feature Aggregation (DFA) module. Experiment results demonstrate that GDGT-OSR can restore more details and richer textures over other methods, achieving SOTA performance on omnidirectional image super-resolution.